# Developmental Symmetry-Loss: A Free-Energy Perspective on Brain-Inspired Invariance Learning

### v1

@@ Abstract
This is a categorical quotient and the functionals $\chi_{1},\ldots,\chi_{m}\in\mathcal{F}^{G}$ form a complete fundamental system of symmetries, meaning that $\mathbf{K}[\chi_{i}:i=1,\ldots,m]=\mathcal{F}^{G}$. Notably, the function $\pi$ serves to identify orbit closures; that is, for any points $v,x\in\Omega$, we have $\overline{Gx}=\overline{Gv}$ if and only if $\pi(x)=\pi(v)$ (kraft1985geometrische; mumford1994geometric).
@@
This extension can be expressed as a short exact sequence of groups,
$1\rightarrow\widehat{G}_{k}\longrightarrow\widehat{G}_{k+1}\longrightarrow\Delta G_{k}\rightarrow 1,\quad\Delta G_{k}:=\widehat{G}_{k+1}/\widehat{G}_{k},$
where $\Delta G_{k}$ measures the quotient action contributed by the newly incorporated symmetry. When $\Delta G_{k}$ acts independently of $\widehat{G}_{k}$ (i.e. their intersection is trivial), the refinement is said to be orthogonal, and $\widehat{G}_{k+1}$ forms a semi–direct product $\widehat{G}_{k+1}\simeq\widehat{G}_{k}\rtimes\Delta G_{k}$. Otherwise, the groups overlap nontrivially, indicating that the newly added symmetry partially reuses or reparametrizes degrees of freedom already represented at earlier layers.
Hence, the refinement hierarchy can be understood as a sequence of nested group extensions,

### v2

@@ Abstract
This is a categorical quotient and the functionals $\chi_{1},\ldots,\chi_{m}\in\mathcal{F}^{G}$ form a complete fundamental system of symmetries, meaning that $\mathbf{K}[\chi_{i}:i=1,\ldots,m]=\mathcal{F}^{G}$. Notably, the function $\pi$ serves to identify orbit closures; that is, for any points $v,x\in\mathcal{X}$, we have $\overline{Gx}=\overline{Gv}$ if and only if $\pi(x)=\pi(v)$ (kraft1985geometrische; mumford1994geometric).
@@
This refinement induces an extension of effective symmetry actions. When $\widehat{G}_{k}$ is normal in $\widehat{G}_{k+1}$, this extension can be represented by a short exact sequence
$1\rightarrow\widehat{G}_{k}\longrightarrow\widehat{G}_{k+1}\longrightarrow\Delta G_{k}\rightarrow 1,\qquad\Delta G_{k}:=\widehat{G}_{k+1}/\widehat{G}_{k},$
where $\Delta G_{k}$ measures the additional symmetry degrees of freedom introduced at refinement step $k$.
In the general case, $\Delta G_{k}$ should be interpreted as an effective quotient capturing new orbit directions modulo the previously internalized symmetries; formally, it may be viewed as a quotient in the category of group actions (equivalently, as a homogeneous space of $\widehat{G}_{k+1}$ modulo $\widehat{G}_{k}$), rather than as an abstract group.
When $\Delta G_{k}$ acts independently of $\widehat{G}_{k}$ (i.e. their intersection is trivial), the refinement is said to be orthogonal, and $\widehat{G}_{k+1}$ forms a semi–direct product $\widehat{G}_{k+1}\simeq\widehat{G}_{k}\rtimes\Delta G_{k}$. Otherwise, the groups overlap nontrivially, indicating that the newly added symmetry partially reuses or reparametrizes degrees of freedom already represented at earlier layers.
Hence, the refinement hierarchy can be understood as a sequence of nested symmetry extensions,